Factoring: Quadratic Equations and completing the square
x-1-6/x=0

Answer by stanbon(75887)

(x-1) - (6/x) = 0
Multiply thru by x to get:
x^2-x-6 = 0
(x-3)(x+2) = 0
x = 3 or x = -2
==================
Cheers,
Stan H.